# Particle Filtering for a Class of State-Space Models with Low and Degenerate Observational Noise

### v1

@@ 4.3 Bridges and Manifolds
where $Y_{n}\in\mathbb{R}^{d_{y}}$, $1\leq d_{y}<d_{x}$, $A_{n}\in\mathbb{R}^{d_{y}\times d_{x}}$ is of full row rank, and independently for each $n\in\mathbb{N}$ $\epsilon_{n}\in\mathbb{R}^{d_{y}}$ are a sequence of independent noises with positive Lebesgue density $p_{n}$ and $\Delta\geq 0$ and $\Sigma_{n}$ is a sequence of real positive definite symmetric matrices. This is the same linear set-up as the previous section. Note that we present the model with observations at regular and unit times only to simplify the notation and make it consistent with previous articles [4]. However, this need not be the case and observations could in principle have a different amount of time in-between them, or indeed be irregularly observed.
4.3.1 Low Noise
Let $\mathbb{W}$ is the law of Brownian motion. The probability measure associated to the smoother in continuous-time and up-to a constant is at time $n$
@@ 4.3.1 Low Noise
In practice we cannot work with this and consider a time discretization.
We first give the standard Euler-Maruyama time discretization of the solution to (4.3) (associated to a time interval [k1,k][k-1,k], $k\in\{1,\dots\}$) on a regular grid of spacing $\Xi_{l}=2^{-l}$, with starting point $x_{k-1}^{\circ}$ and ending point $x_{k}^{\circ}$. The choice $\Xi_{l}=2^{-l}$, is again for consistency of presentation with [4], but other schemes can be used in practice. We have, for $j\in\{0,1\dots,\Xi_{l}^{-1}-2\}$:
$X_{k-1+(j+1)\Delta_{l}}^{\circ}$ $=X_{k-1+j\Xi_{l}}^{\circ}+\mu_{k}^{\circ}(k-1+j\Xi_{l},X^{\circ}_{k-1+j\Xi_{l}};x_{k}^{\circ})\Xi_{l}+$
$\sigma(X^{\circ}_{k-1+j\Xi_{l}})\left[W_{k-1+(j+1)\Xi_{l}}-W_{k-1+j\Xi_{l}}\right].$ (4.7)
@@
As a result of the time-discretization we focus on the smoother (up-to a normalizing constant):
@@
We set
@@
  • Iterate: For $i\in\{1,\dots,N\}$, sample $\widetilde{a}_{n-1}^{i}|\widetilde{z}_{n-2}^{1:N},\mathbf{w}_{[n-2,n-1]}^{1:N,l},\widetilde{z}_{n-1}^{1:N}$ i.i.d. from
    $\frac{w_{n-1}^{\Delta,l}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{i}},\mathbf{w}_{[n-2,n-1]}^{\widetilde{a}_{n-2}^{i},l},\widetilde{z}_{n-1}^{\widetilde{a}_{n-1}^{i}})}{\sum_{s=1}^{N}w_{n-1}^{\Delta,l}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{s}},\mathbf{w}_{[n-2,n-1]}^{\widetilde{a}_{n-2}^{s},l},\widetilde{z}_{n-1}^{s})}.$
    For $i\in\{1,\dots,N\}$, sample $\widetilde{Z}_{n}^{i}|\widetilde{z}_{n-1}^{\widetilde{a}_{n-1}^{i}}$ independently from $\widetilde{q}_{n}^{\Delta}(\cdot|\widetilde{z}_{n-1}^{\widetilde{a}_{n-1}^{i}})$ and independently $\mathbf{w}_{[n-1,n]}^{i,l}\sim\mathbb{W}(\cdot)$. Set $n=n+1$ and return to the start of the bullet point.
@@ 4.3.2 Degenerate Noise
We set
$w_{n}^{\star,l}(z_{n-1},\mathbf{w}_{[n-1,n]}^{l},z_{n})$ $=R_{n-1,n}^{l}(C_{n-1,n}^{l}(u_{n-1}^{\star}(z_{n-1}),\mathbf{w}_{[n-1,n]}^{l},u_{n}^{\star}(z_{n})))\frac{1}{q_{n}^{\star}(z_{n}|z_{n-1})}.$
@@
  • Iterate: For $i\in\{1,\dots,N\}$, sample $a_{n-1}^{i}|z_{n-2}^{1:N},\mathbf{w}_{[n-2,n-1]}^{1:N,l},z_{n-1}^{1:N}$ i.i.d. from
    $\frac{w_{n-1}^{\star,l}(z_{n-2}^{a_{n-2}^{i}},\mathbf{w}_{[n-2,n-1]}^{a_{n-2}^{i}},z_{n-1}^{a_{n-1}^{i}})}{\sum_{s=1}^{N}w_{n-1}^{\star,l}(z_{n-2}^{a_{n-2}^{s}},\mathbf{w}_{[n-2,n-1]}^{a_{n-2}^{s}},z_{n-1}^{s})}.$
    For $i\in\{1,\dots,N\}$, sample $Z_{n}^{i}|z_{n-1}^{a_{n-1}^{i}}$ independently from $q_{n}^{\star}(\cdot|z_{n-1}^{a_{n-1}^{i}})$ and independently $\mathbf{w}_{[n-1,n]}^{i,l}\sim\mathbb{W}(\cdot)$. Set $n=n+1$ and return to the start of the bullet point.
The algorithm here is presumably the uniformly elliptic embodyment of the ideas in [6].
@@ 5.1 Linear Gaussian Model
$X_{n}=B_{n}X_{n-1}+\nu_{n}$ (5.2)
where, independently for each $n\in\mathbb{N}$, $\nu_{n}\sim\mathcal{N}_{d_{x}}(0,\Omega)$. We choose $X_{0}=(0,0)^{\top}$, $\Omega=\mathtt{I}_{d_{x}}$, $B_{n}=0.9\mathtt{I}_{d_{x}}$, $(A_{n})_{ij}=1/d_{x}$ (computes the average of the state), $n\in\{1,\dots,20\}$ and $\Delta=10^{-4}$. We run the Kalman filter, the particle filters with bootstrap and optimal proposals under the natural parameterization and the proposed filter under the noise parameterization. We set the number of particles to $N=10^{4}$ and resample if the effective sample size (ESS) drops below $N/2$. In Figure 1, we plot the marginal densities of the first component of $X_{n}$ at various $n$. We see that, except the bootstrap, all filters yield similar estimates as one might expect in this case.
@@
We now repeat the experiment for various $\Delta$ and alllow $\Delta\to 0$. We monitor the median ESS (Figure 2) across the time-steps for a given $\Delta$. In Figure 2 we observe that the bootstrap PF degenerates for relatively large $\Delta$, whilst the other methods remain stable. The optimal proposal under natural parameterization becomes unstable for very tiny $\Delta$, due to numerical errors and the reasoning in Remark 3.3. We can also see it in the mean squared error (MSE) of the mean estimates (error averaged across time), where the bootstrap PF collapses but the low-noise optimal proposal converges to the degenerate case.
@@ 5.2 Nonlinear Model with Additive Noise
Now we consider a nonlinear Lorenz-96 model with additive Gaussian noise, $d_{x}=8$
$X_{n}=F(X_{n-1})+\nu_{n}$ (5.3)
where $F_{i}(x)=x_{i}+\Delta t((x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F_{0})$, $i\in\{1,\dots,d_{x}\}$ and, independently for each $n\in\mathbb{N}$, $\nu_{n}\sim\mathcal{N}_{d_{x}}(0,\Omega)$. We fix the parameter of the model to $F_{0}=8$ to make the dynamics chaotic, and simulate it at stepsize $\Delta t=10^{-2}$ for $n\in\{1,\dots,400\}$ to obtain the data, starting from $X_{0}=(1,1)^{\top}$. We fix $\Delta=10^{-4}$ and $\Omega=\Delta t\mathtt{I}_{d_{x}}$, the matrix $A_{n}$ is constructed such that we observe the $1^{\text{s}t}$ and $5^{\text{t}h}$ component of $X_{n}$. We can observe the chaotic behavior of the model in Figure 3, which are the simulated dynamics of the model.
@@
We run all the mentioned filters (except for Kalman filter which does not apply here) again for $N=10^{4}$ particles with the same resampling rule. We plot several marginals in Figure 4, where we recover similar results as in the case of the linear Gaussian model, except for the case of the bootstrap PF; the algoritm performs so badly that the plots are essentially useless. Further, we plot the ESS versus the time index in Figure 5, where we see that the bootstrap PF collapses from the very beginning, while other methods give adequate results in terms of this measure. To further examine the accuracy of the PFs we observe individual sample paths of the partices (for a given run) and their mean for $8^{\text{t}h}$ component in Figure 6. In both the top and bottom of Figure 6 the true value is the actual signal used to generate the data. Apart from the bootstrap PF all of the methods are seemingly able to recover the true hidden state in the $8^{\text{t}h}$ component. In Figure 7 we consider re-running all of the methods for different $\Delta$ and sending $\Delta\to 0$. Figure 7 displays the median ESS (accross time) of the four PFs. In this case the optimal proposal under the natural parameterization does not seem to collapse (in terms of ESS) but this could be as the weight does not degenerate to zero.
@@
Figure 7: ESS versus $\log_{10}(1/\Delta)$. This is for the nonlinear model.
@@ 5.3 Diffusion Model
$dX_{t}=\mu(X_{t})dt+\sigma(X_{t})dW_{t}$ (5.4)
with $\mu(x)=((x_{1}-x_{1}^{3}-x_{2})/\alpha,\gamma x_{1}-x_{2}+\beta)^{\top}$ and $\sigma(x)=\sigma_{0}\mathtt{I}_{d_{x}}$. This is an elliptic diffusion with constant diffusion term. We fix the model parameters to $\alpha=0.1$, $\gamma=1$, $\beta=0.2$ and $\sigma_{0}=0.1$. We choose the starting values to be $X_{0}=(0.5,0.5)^{\top}$. We simulate the dynamics in the interval $t\in[0,10]$ using the Euler-Maruyama scheme with stepsize $\Xi=5\times 10^{-3}$ and observe the first component at regular times with spacing $10^{-2}$ and finally $\Delta=10^{-8}$ is the noise level. This scenario leads to 20 discretization points between two consecutive observations and 2000 in total.
We consider an auxiliary process with linearized drift $\tilde{\mu}(x)=\mu(x_{\text{ref}})+\frac{\partial}{\partial x}\mu(x_{\text{ref}})(x-x_{\text{ref}})$. We can calculate the reference point by evolving the true dynamics deterministically starting from $x$ using the stepsize $\Xi$ for 20 iterations until we reach the next observation time. The auxiliary process is an Ornstein-Uhlenbeck process with a transition density of $x^{\prime}$ given $x$ over time step $h$ as $\mathcal{N}_{2}(m,Q)$ where
$m=\exp(Jh)x+J^{-1}(\exp(Jh)-I)(\mu-Jx_{\text{ref}})$
@@
The bootstrap PF will use a proposal that is exactly the auxiliary process (on $x_{k}$ the Brownian motion is sampled from its law) whereas the optimal (which is not optimal in any sense) is a proposal proportional to the term $h_{k}(y_{k}|x_{k})$ multiplied by the density of the auxiliary process and the Brownian motion sampled from its law.
In Figure 8 we plot the ESS against time for a particular run. Here, we have set $N=10^{4}$. As in the previous example, the bootstrap PF collapses completely and the other filters perform quite well for this criterion. This shows that one can filter in low or degenerate noise and with a high amount of time discretization. This deduction is supported by the results in Figures 9 (marginal densities of the state) and 10 (tracking the true signal) which shows that one can reconstruct the signal process in this case. In Figure 11 we repeat the experiment for $\Delta\to 0$, where we see how the bootstrap PF collapses and the low-noise PF converges to the degenerate case.
Figure 8: ESS versus time. This is for the diffusion model.
Figure 9: Marginals of $X_{t,2}$ for $t\in\{0.1,3.4,6.7,10\}$. This is for the diffusion model.
@@
Figure 11: ESS versus $\log_{10}(1/\Delta)$. This is for the diffusion model.

### v2

@@ 4.3 Bridges and Manifolds
where $Y_{n}\in\mathbb{R}^{d_{y}}$, $1\leq d_{y}<d_{x}$, $A_{n}\in\mathbb{R}^{d_{y}\times d_{x}}$ is of full row rank, and independently for each $n\in\mathbb{N}$ $\epsilon_{n}\in\mathbb{R}^{d_{y}}$ are a sequence of independent noises with positive Lebesgue density $p_{n}$ and $\Delta\geq 0$. This is the same linear set-up as the previous section. Note that we are considering hidden state that evolves in continuous time $t$ that is observed at uniformly spaced discrete times and in particular when $t=n=1,2,3,\ldots$, i.e. at unit time intervals. This allows to simplify the notation and make it consistent with previous articles [4]. However, this need not be the case and observations could in principle have a different amount of time in-between them, or indeed be irregularly observed.
4.3.1 The Low Noise Case
Let $\mathbb{W}$ is the law of Brownian motion. The probability measure associated to the smoother in continuous-time at time $n$ will be proportional to
@@ 4.3.1 The Low Noise Case
In practice we cannot work with this and consider a time discretization. We first give the standard Euler-Maruyama time discretization of the solution to (4.3) (associated to a time interval [k1,k][k-1,k], $k\in\{1,\dots\}$) on a regular grid of spacing $\Xi_{l}=2^{-l}$, with starting point $x_{k-1}^{\circ}$ and ending point $x_{k}^{\circ}$. The choice $\Xi_{l}=2^{-l}$, is again for consistency of presentation with [4], but other schemes can be used in practice. We have, for $j\in\{0,1\dots,\Xi_{l}^{-1}-2\}$:
$X_{k-1+(j+1)\Xi_{l}}^{\circ}$ $=X_{k-1+j\Xi_{l}}^{\circ}+\mu_{k}^{\circ}(k-1+j\Xi_{l},X^{\circ}_{k-1+j\Xi_{l}};x_{k}^{\circ})\Xi_{l}+$
$\qquad\qquad\qquad\sigma(X^{\circ}_{k-1+j\Xi_{l}})\left[W_{k-1+(j+1)\Xi_{l}}-W_{k-1+j\Xi_{l}}\right].$ (4.7)
@@
As a result of the time-discretization we can target the smoother that is proportional to
@@
For the importance sampling part, we proceed similar to Section 3.1.1 and for the weights set:
@@
  • Iterate: For $i\in\{1,\dots,N\}$, sample $\widetilde{a}_{n-1}^{i}\sim\text{Cat}(\cdot;N,\mathtt{w}^{\Delta}_{n})$ with
    $\mathtt{w}^{\Delta}_{n}=\left(\frac{w_{n-1}^{\Delta,l}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{1}},\mathbf{w}_{[n-2,n-1]}^{\widetilde{a}_{n-2}^{1},l},\widetilde{z}_{n-1}^{1})}{\sum_{s=1}^{N}w_{n-1}^{\Delta,l}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{s}},\mathbf{w}_{[n-2,n-1]}^{\widetilde{a}_{n-2}^{s},l},\widetilde{z}_{n-1}^{s})},\ldots,\frac{w_{n-1}^{\Delta,l}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{N}},\mathbf{w}_{[n-2,n-1]}^{\widetilde{a}_{n-2}^{N},l},\widetilde{z}_{n-1}^{N})}{\sum_{s=1}^{N}w_{n-1}^{\Delta,l}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{s}},\mathbf{w}_{[n-2,n-1]}^{\widetilde{a}_{n-2}^{s},l},\widetilde{z}_{n-1}^{s})}\right)$
    For $i\in\{1,\dots,N\}$, sample independently $\widetilde{Z}_{n}^{i}|\widetilde{z}_{n-1}^{\widetilde{a}_{n-1}^{i}}$ from $\widetilde{q}_{n}^{\Delta}(\cdot|\widetilde{z}_{n-1}^{\widetilde{a}_{n-1}^{i}})$ and $\mathbf{w}_{[n-1,n]}^{i,l}\sim\mathbb{W}(\cdot)$. Set $n=n+1$ and return to the start of the bullet point.
@@ 4.3.2 Degenerate Noise
Similar to Section 3.1.2, for the importance weights we set:
$w_{n}^{\star,l}(z_{n-1},\mathbf{w}_{[n-1,n]}^{l},z_{n})$ $=\frac{R_{n-1,n}^{l}(C_{n-1,n}^{l}(u_{n-1}^{\star}(z_{n-1}),\mathbf{w}_{[n-1,n]}^{l},u_{n}^{\star}(z_{n})))}{q_{n}^{\star}(z_{n}|z_{n-1})}.$
@@
  • Iterate: For $i\in\{1,\dots,N\}$, Sample $a_{n-1}^{i}\sim\text{Cat}(\cdot;N,\mathtt{w}^{\star,l}_{n-1})$ i.i.d. with
    $\mathtt{w}^{\star,l}_{n-1}=\left(\frac{w_{n-1}^{\star,l}(z_{n-2}^{a_{n-2}^{1}},\mathbf{w}_{[n-2,n-1]}^{a_{n-2}^{1}},z_{n-1}^{1})}{\sum_{s=1}^{N}w_{n-1}^{\star,l}(z_{n-2}^{a_{n-2}^{s}},\mathbf{w}_{[n-2,n-1]}^{a_{n-2}^{s}},z_{n-1}^{s})},\ldots,\frac{w_{n-1}^{\star,l}(z_{n-2}^{a_{n-2}^{N}},\mathbf{w}_{[n-2,n-1]}^{a_{n-2}^{N}},z_{n-1}^{N})}{\sum_{s=1}^{N}w_{n-1}^{\star,l}(z_{n-2}^{a_{n-2}^{s}},\mathbf{w}_{[n-2,n-1]}^{a_{n-2}^{s}},z_{n-1}^{s})}\right).$
    For $i\in\{1,\dots,N\}$, sample independently $Z_{n}^{i}|z_{n-1}^{a_{n-1}^{i}}$ from $q_{n}^{\star}(\cdot|z_{n-1}^{a_{n-1}^{i}})$ and $\mathbf{w}_{[n-1,n]}^{i,l}\sim\mathbb{W}(\cdot)$. Set $n=n+1$ and return to the start of the bullet point.
The algorithm here is presumably a PF implementation based on the ideas in [7] for the uniformly elliptic case.
@@ 5.1 Linear Gaussian Model
$X_{n}=B_{n}X_{n-1}+\nu_{n}$
where, independently for each $n\in\mathbb{N}$, $\nu_{n}\sim\mathcal{N}_{d_{x}}(0,\Omega)$. We choose $X_{0}=(0,\dots,0)^{\top}$, $\Omega=\mathtt{I}_{d_{x}}$, $B_{n}=0.9\mathtt{I}_{d_{x}}$, $(A_{n})_{ij}=1/d_{x}$ (computes the average of the state), $n\in\{1,\dots,20\}$ and $\Delta=10^{-4}$. We run the Kalman filter, the particle filters with bootstrap and optimal proposals under the natural parameterization and the proposed filter under the noise parameterization. We set the number of particles to $N=10^{4}$ and resample if the effective sample size (ESS) drops below $N/2$. In Figure 1, we plot the marginal densities of the first component of $X_{n}$ at various $n$. We see that, except the bootstrap, all filters yield similar estimates as one might expect in this case.
@@
We now repeat the experiment for various $\Delta$ and alllow $\Delta\to 0$. We monitor the median ESS (Figure 2) across the time-steps for a given $\Delta$. In Figure 2 we observe that the bootstrap PF degenerates for relatively large $\Delta$, whilst the other methods remain stable. We can also see it in the mean squared error (MSE) of the mean estimates (error averaged across time and dimensions, see Equation 5.1), where the bootstrap PF collapses but the low-noise optimal proposal converges to the degenerate case.
$\text{MSE}=\frac{1}{n_{\text{step}}}\sum_{n=1}^{n_{\text{step}}}\left\{\frac{1}{d_{x}}\sum_{j=1}^{d_{x}}\left[\left(\frac{1}{N}\sum_{i=1}^{N}X_{n,j}^{i}\right)-\mathbb{E}\left(X_{n,j}\right)\right]^{2}\right\}$ (5.1)
@@ 5.2 Nonlinear Model with Additive Noise
Now we consider a nonlinear Lorenz-96 model with additive $t$-distributed noise, $d_{x}=8$
$X_{n}=F(X_{n-1})+\nu_{n}$
where $F_{j}(x)=x_{j}+\Delta t((x_{j+1}-x_{j-2})x_{j-1}-x_{j}+F_{0})$, $j\in\{1,\dots,d_{x}\}$ and, independently for each $n\in\mathbb{N}$, $\nu_{n}\sim t_{d_{x}}(0,\Omega,\vartheta)$ (multivariate $t$-distribution with scale matrix $\Omega$ and $\vartheta>0$ degrees of freedom). We fix the parameter of the model to $F_{0}=8$ and the degrees of freedom of the innovation noise to $\vartheta=2$, and simulate it at stepsize $\Delta t=10^{-2}$ for $n\in\{1,\dots,400\}$ to obtain the data, starting from $X_{0}=(1,\dots,1)^{\top}$. We fix $\Delta=10^{-4}$ and $\Omega=\Delta t\mathtt{I}_{d_{x}}$, the matrix $A_{n}$ is constructed such that we observe the $1^{\text{s}t}$ and $5^{\text{t}h}$ component of $X_{n}$. We can observe the chaotic behavior of the model in Figure 3, which are the simulated dynamics of the model.
@@
Note that the optimal proposal is not available directly due to $\nu_{n}$ being $t$-distributed. However, it admits a hierarchical representation as a scale-mixture of Gaussians, via $\nu_{n}|\omega_{n}\sim\mathcal{N}(0,\Omega/\omega_{n})$ with a Gamma-distributed auxiliary variable $\omega_{n}\sim\Gamma(\vartheta/2,\vartheta/2)$. Given $\omega_{n}$ the noise is conditionally Gaussian, so can use the optimal proposals from the Gaussian case to obtain a conditionally optimal proposal for both the natural and noise parameterizations, conditioned on $\omega_{n}$, which we sample from the prior $\Gamma(\vartheta/2,\vartheta/2)$.
We run all the mentioned filters (except for Kalman filter which does not apply here) again for $N=10^{4}$ particles with the same resampling rule. We plot several marginals in Figure 4, where we recover similar results as in the case of the linear Gaussian model, except for the case of the bootstrap PF whose performance is too poor to give presentable results. Further, we plot the ESS versus the time index in Figure 5, where we see that the bootstrap PF collapses from the very beginning, while other methods give adequate results in terms of this measure. To further examine the accuracy of the PFs we plot individual sample paths of the particles (for a given run) and their mean for $8^{\text{t}h}$ component in Figure 6. In both the top and bottom of Figure 6 the true value is the actual signal used to generate the data. Apart from the bootstrap PF all of the methods are seemingly able to recover the true hidden state in the $8^{\text{t}h}$ component. In Figure 7 we consider re-running all of the methods for different $\Delta$ and sending $\Delta\to 0$. Figure 7 displays the median ESS (across time) and the MSE (between estimated mean and the true signal) of the four PFs. In this case the optimal proposal under the natural parameterization does not seem to collapse (in terms of ESS) but this could be as the weight does not degenerate to zero. The MSE of the bootstrap is too large as it collapses immediately. The other two seem to roughly approach the MSE of the degenerate PF, but there is still variation due to iterates not being exactly the same.
@@
Figure 7: ESS (left) and MSE (right) versus $\log_{10}(1/\Delta)$. This is for the nonlinear model.
@@ 5.3 Diffusion Model
$dX_{t}=\mu(X_{t})dt+\sigma(X_{t})dW_{t}$
with $\mu(x)=((x_{1}-x_{1}^{3}-x_{2})/\alpha,\gamma x_{1}-x_{2}+\beta)^{\top}$ and $\sigma(x)=\sigma_{0}\mathtt{I}_{d_{x}}$. This is an elliptic diffusion with constant diffusion term. We fix the model parameters to $\alpha=0.1$, $\gamma=1$, $\beta=0.2$ and $\sigma_{0}=0.1$. We choose the starting values to be $X_{0}=(0.5,0.5)^{\top}$. We simulate the dynamics in the interval $t\in[0,10]$ using the Euler-Maruyama scheme with stepsize $\Xi=5\times 10^{-3}$ and observe the first component at regular times with spacing $10^{-2}$ and finally $\Delta=10^{-8}$ is the noise level. This scenario leads to 20 discretization points between two consecutive observations and 2000 in total. The simulated dynamics is given in Figure 8.
Figure 8: Components of $X_{t}$ for $t\in[0,10]$. This is for the diffusion model.
We consider an auxiliary process with linearized drift $\tilde{\mu}(x)=\mu(x_{\text{ref}})+\frac{\partial}{\partial x}\mu(x_{\text{ref}})(x-x_{\text{ref}})$. We can calculate the reference point $x_{\text{ref}}$ by evolving the true dynamics deterministically starting from $x$ using the stepsize $\Xi$ for 20 iterations until we reach the next observation time. The auxiliary process is an Ornstein-Uhlenbeck process with a transition density of $x^{\prime}$ given $x$ over time step $h$ as $\mathcal{N}_{2}(m,Q)$ where
$m=\exp(Jh)x+J^{-1}(\exp(Jh)-I)(\mu-Jx_{\text{ref}})$ (5.2)
@@
The bootstrap PF will use a proposal that is exactly the auxiliary process (on $x_{k}$ the Brownian motion is sampled from its law) whereas the optimal (which is not optimal in any sense) is a proposal proportional to the term $h_{k}(y_{k}|x_{k})$ multiplied by the density of the auxiliary process given by Equation 5.2 and the Brownian motion sampled from its law.
In Figure 9 we plot the ESS against time for a particular run. Here, we have set $N=10^{4}$ and resample at each step. As in the previous example, the bootstrap PF collapses completely and the other filters perform quite well for this criterion. This shows that one can filter in low or degenerate noise and with a very fine time discretization. This deduction is supported by the results in Figures 10 (marginal densities of the state) and 11 (tracking the true signal) which shows that one can reconstruct the signal process in this case. In Figure 12 we repeat the experiment for $\Delta\to 0$, where we see how the bootstrap PF collapses and the low-noise PF converges to the degenerate case.
Figure 9: ESS versus $n$. This is for the diffusion model.
Figure 10: Marginals of $X_{t,2}$ for $n\in\{1,34,67,100\}$. This is for the diffusion model.
@@
Figure 12: ESS (left) and MSE (right) versus $\log_{10}(1/\Delta)$. This is for the diffusion model.